# Joint State and Noise Covariance Estimation

### v1

@@ IV-A Inner Subproblem: Covariance Estimation
$\underset{x\in\mathcal{M}}{\text{minimize}}&\qquad\underset{P\succeq 0}{\min}&&\,\,F(x,P)\\ &\quad\quad\text{subject to}&&\text{appropriate constraints on $P$},\end{aligned}$ (9)
@@ IV-B Analytical Solution to the Inner Subproblem
$&\underset{P\succeq 0}{\text{minimize}}&&-\log\det P+\langle M(x),P\rangle\\ &\text{subject to}&&\text{appropriate constraints on $P$}.\end{aligned}$ (10)
@@
Remark 4 (ML Estimation).
As we mentioned earlier, the objective function in the ML formulation for estimating the noise covariance can be written as $-\log\det P+\langle S(x),P\rangle$ which is similar to (10). However, unlike $M(x)$, the sample covariance $S(x)$ can become singular. Therefore, Theorem 1 readilly applies to the ML case (and its constrained variants) with an important exception: without the constraint $\Sigma\succeq\lambda_{\text{min}}I$, if $S(x)$ becomes singular, the problem becomes unbounded from below and thus ML estimation becomes ill-posed.
Remark 5 (Singularities of $S(x)$).
According to Theorem 1, when $S(x)$ is singular, the spectrum of $M(x)$ and consequently solution to the inner subproblems associated with Problems 1 and 2 become more influenced by the Wishart prior than the observed data. The sample covariance matrix could be singular in many different situations, but two particular cases that can lead to singularity are as follows:
@@
Theorem 1 has a clear geometric interpretation. For instance, the optimal covariance matrix $\Sigma^{\star}(x)={P^{\star}(x)}^{-1}$ for Problem 3 (14) has the following properties: (i) it preserves the eigenvectors of $M(x)$, which define the principal axes of the associated confidence ellipsoid; (ii) it matches $M(x)$ along directions corresponding to eigenvalues within the range $[\lambda_{\text{min}},\lambda_{\text{max}}]$; and (iii) it only adjusts the radii of the ellipsoid along the remaining principal axes to satisfy the constraint. This is visualized in Figure 2. Similarly, the confidence ellipsoid associated to (12) is obtained by projecting the confidence ellipsoid of $M(x)$ onto the standard basis. In the case of (16), the radii of the projected ellipsoid are adjusted to satisfy the eigenvalue constraint.
@@ IV-C2 Preprocessed Measurements and Non-Additive Noise
In SLAM, raw measurements are often preprocessed and transformed nonlinearly into standard models supported by popular solvers. For instance, raw range-bearing measurements (corrupted by additive noise with covariance $\Sigma_{\text{true}}$) are often expressed in Cartesian coordinates. As a result, although the raw measurements generated by the sensor have identically distributed noise, the transformed measurements that appear in the least squares problem may have different covariances because of the nonlinear transformation. Let $\Sigma_{\text{true}}$ be the covariance of raw measurements, and $\Sigma_{i}$ be the covariance of the $i$th transformed measurement. In practice, $\Sigma_{i}$ is approximated by linearization, i.e., $\Sigma_{i}\approx J_{i}\Sigma_{\text{true}}J_{i}^{\top}$ in which $J_{i}$ is the (known) Jacobian of the transformation. It is easy to verify that Theorem 1 readilly extends to this case when $J_{i}$’s are full-rank square matrices by replacing the sample covariance $S(x)$ as defined in (3) with
@@ V-B Block-Coordinate Descent
    Fix $P$ to its most recent value and minimize the joint MAP objective function in Problems 1-4 with respect to $x\in\mathcal{M}$. This results in a standard nonlinear least squares problem (where residuals are weighted by $P$) over $\mathcal{M}$, which can be (locally) solved using existing solvers such as [2, 10, 17].
@@
We explore two variants of this procedure shown in Algorithms 3 and 4. The BCD algorithms of the type considered here address the limitations of Algorithm 2. Specifically, the problem in the first step can be readily solved using standard solvers widely used in robotics and computer vision, which are also capable of exploiting sparsity in residuals. The second step is highly efficient, as it only requires computing the $m\times m$ matrix $M(x)$ as defined in (5), which can be done in $O(k\cdot m^{2})$ time. In the case of Problem 3, one must also compute the eigendecomposition of $M(x)$. In practice, $m$ (the dimension of the residuals) is typically a small constant (i.e., $m=O(1)$; e.g., in 3D PGO, $m=6$), and therefore the overall time complexity of the second step of BCD is $O(k)$, i.e., linear in the number of measurements.
@@
2:  $t\leftarrow 1$
@@
4:  $P^{t}\leftarrow P^{\star}(x_{\text{init}})$
5:  while $t\leq\tau$ do
6:    // Step 1: Update $x^{t}$ using a descent step using retraction $\text{Retr}_{x^{t}}(\cdot)$, Riemannian gradient $\mathrm{grad}_{x}\,R(\cdot)$ with respect to $x$, and step-size $\eta$; see Theorem 2 and Appendix E.
7:   $x^{t}\leftarrow\text{Retr}_{x^{t}}(-\eta\,\mathrm{grad}_{x}\,R(x^{t},P^{t}))$
8:    // Step 2: optimize $P$ (Theorem 1)
9:   $P^{t}\leftarrow P^{\star}(x^{t})$
10:   $t\leftarrow t+1$
@@
Algorithm 4 Block-Exact BCD for Joint MAP
1:procedure BCD($x_{\text{init}}$)
2:  $t\leftarrow 1$
3:   // Initialize the information matrix
4:  $P^{t}\leftarrow P^{\star}(x_{\text{init}})$
5:  while not converged do
6:    // Step 1: optimize $x$
7:   $x^{t}\in\operatorname*{argmin}_{x\in\mathcal{M}}\,\,\frac{1}{2}\sum_{i=1}^{k}\|r_{i}(x)\|^{2}_{P^{t}}$
8:    // Step 2: optimize $P$ (Theorem 1)
9:   $P^{t}\leftarrow P^{\star}(x^{t})$
10:   $t\leftarrow t+1$
11:  end while
12:  return $(x^{t},P^{t})$
13:end procedure
Before studying the convergence properties of these algorithms we introduce the necessary assumption below. See Appendix E for background information. For simplifying the presentation we define $R(x,P)\triangleq\langle M(x),P\rangle$.
@@
See Appendix D. ∎
As can be seen in Algorithm 4, one might be able to solve the optimization problems associated with each of the coordinates uniquely and exactly. For example, for the case where the residuals $r_{i}(x)$ are affine functions of $x$ and $\mathcal{M}$ is convex, the optimization problem associated with $x$ has a unique solution (assuming a non-singular Hessian) can be solved exactly. This case satisfies the following assumption.
@@
The proof of the theorem follows directly from [20, Theorem 1]. ∎
@@ VI-A Linear Measurement Model
in which $\Sigma^{\text{base}}\succeq 0$ is a fixed random covariance matrix, and $\sigma^{2}\in\{10^{-2},10^{-1},1,10,10^{2}\}$ is a variable that controls the noise level in our experiments. We did not impose a prior on $P$ (i.e., prior weight $w_{\text{prior}}=0$; see (31) in Appendix A), leading to unconstrained joint ML estimation of $x_{\text{true}}$ and $\Sigma_{\text{true}}$.
@@ VI-B Pose-Graph Optimization Ablations
Dataset: We used a popular synthetic PGO benchmark, the Manhattan dataset [19], and generated new measurement realizations with varying values of the actual noise covariance matrix. The dataset consists of $3,500$ poses and $k=5,598$ relative-pose measurements. This dataset is notoriously poorly connected [16]. Therefore, to analyze the effect of connectivity on covariance estimation, we also performed experiments on modified versions of this dataset, where additional loop closures were introduced by connecting pose $i$ to poses $i+2$ and $i+3$. This modification increased the total number of measurements to $k=12,593$.
@@
Algorithms: We implemented a variant of Algorithm 3 in C++, based on g2o [17], for PGO problems. Instead of using Riemannian gradient descent, we used g2o’s implementation of Powell’s Dog-Leg method [21] on $\mathrm{SE}(2)$ to optimize $x$. In each outer iteration, our implementation retrieves the residuals for each measurement, computes $M(x)$ as defined in (5), and updates the noise covariance estimate using Theorem 1. To handle cases where $w_{\text{prior}}=0$ (i.e., ML estimation) and $S(x)$ is poorly conditioned, we set $\lambda_{\text{min}}=10^{-4}$ and $\lambda_{\text{max}}=10^{4}$ in all experiments. We then perform a single iteration of Powell’s Dog-Leg method [21] to update the primary parameters $x$, initializing the solver at the latest estimate of $x$. To ensure convergence across all Monte Carlo trials, we ran 13 outer iterations.
@@
    BCD: In line 9 of Algorithm 3, the noise information matrix is estimated using the ML estimate (i.e., $w_{\text{prior}}\to 0$) subject to eigenvalue constraints. This is equivalent to (14) after replacing $M(x)$ with the sample covariance $S(x)$ at the current value for $x$.
@@
    BCD (diag): In line 9 of Algorithm 3, the noise information matrix is estimated using the ML estimate (as above), subject to both diagonal and eigenvalue constraints. This is equivalent to (16) after replacing $M(x)$ with the sample covariance $S(x)$ at the current value for $x$.
@@
    BCD (Wishart): In line 9 of Algorithm 3, the information matrix is updated according to (14); i.e., using the MAP estimate with a Wishart prior and under the eigenvalue constraints.
@@
    BCD (diag+Wishart): In line 9 of Algorithm 3, the noise information matrix is updated according to (16); i.e., using the MAP estimate with a Wishart prior and under the diagonal and eigenvalue constraints.
@@
[…]
@@ VI-C RIM Dataset
In this section, we present qualitative results on RIM, a large real-world 3D PGO dataset collected at Georgia Tech [8]. This dataset consists of $10,195$ poses and $29,743$ measurements. The g2o dataset includes a default noise covariance matrix, but with these default values, g2o struggles to solve the problem. Specifically, the Gauss-Newton and Dog-Leg solvers terminate after a single iteration, while the Levenberg-Marquardt solver completes 10 iterations but produces the trajectory estimate shown in Figure 6(a).
@@ VII Limitations
This paper did not explore the issue of outliers. Our BCD-type methods can be viewed as (unconventional) instances of iteratively reweighted least squares (IRLS) techniques, where we estimate the entire weight matrix instead of a scalar weight. Consequently, we believe that our joint MAP estimation framework can be easily adapted to integrate with M-estimation IRLS techniques commonly used for outlier-robust estimation, to accurately estimate both the primary parameters and the noise covariance matrix (for inlier measurements) when the dataset contains outliers. We are currently exploring this direction and also aim to implement our method in real-time visual and LiDAR-based SLAM and odometry systems such as [7, 23].
Moreover, our Theorem 2 requires $\mathcal{M}$ to be compact. This result therefore does not immediately apply to problems where $x$ contains translational components. However, we believe we can address this issue in SLAM by confining translational components to potentially large but bounded subsets of $\mathbb{R}^{d}$ (where $d\in\{2,3\}$ is the problem dimension).
Finally, we acknowledge the extensive research on IGLS and Feasible GLS (FGLS) conducted as early as 1970s in econometrics and statistics, which remain largely unknown in engineering; see relevant references in [22, Chapter 12.5] and [26]. Exploring this rich literature may result in new insights and methods that can further improve noise covariance estimation in SLAM and computer vision applications.

### v3

@@ IV-A Inner Subproblem: Covariance Estimation
$\begin{aligned} \underset{x\in\mathcal{M}}{\text{minimize}}&\qquad\underset{P\succeq 0}{\min}&&\,\,F(x,P)\\ &\quad\quad\text{subject to}&&\text{appropriate constraints on $P$},\end{aligned}$ (13)
@@ IV-B Analytical Solution to the Inner Subproblem
$\begin{aligned} &\underset{P\succeq 0}{\text{minimize}}&&-\log\det P+\langle M(x),P\rangle\\ &\text{subject to}&&\text{appropriate constraints on $P$}.\end{aligned}$ (14)
@@
The following theorem provides analytical expressions for the unique optimal solutions to the inner subproblems (14) in Problems 1-4.
@@
As we saw in Remark 1, the objective function in the ML formulation for estimating the noise covariance can be written as $\log\det P+\langle S(x),P\rangle$ which has a similar form to the cost function in (14). However, unlike $M(x)$, the sample covariance $S(x)$ can become singular. Therefore, Theorem 1 readily applies to the ML case (and its constrained variants) with an important exception: without the constraint $\Sigma\succeq\lambda_{\text{min}}I$, if $S(x)$ becomes singular, the problem becomes unbounded from below and thus ML estimation becomes ill-posed. This is formally proved in the following theorem.
Theorem 2.
Consider the (unconstrained) joint ML estimation problem (Remark 1),
$\underset{x\in\mathcal{M},P\succeq 0}{\text{minimize}}\,\,-\log\det P+\left\langle S(x),P\right\rangle.$ (21)
If $S(x)$ is singular, this problem (and the corresponding “inner subproblem”) is unbounded below and thus does not have a solution.
Proof.
See Appendix D. ∎
Remark 4.
Theorem 2 shows that, without the Wishart prior, if $S(x)$ is singular, the unconstrained joint estimation problem (specifically, without $\Sigma\succeq\lambda_{\text{min}}I$) becomes ill-posed. Similarly, singularity of $S(x)$ can make Problems 1 and 2 ill-conditioned when the prior is weak (i.e., $w_{\text{prior}}$ is small) as the corresponding objective function becomes very sensitive in certain directions that correspond to the smallest eigenvalues of $M(x)$. The sample covariance matrix $S(x)$ could be singular in many different situations, but two particular cases that can lead to singularity are as follows:
@@
Theorem 1 has a clear geometric interpretation. For instance, the optimal covariance matrix $\Sigma^{\star}(x)={P^{\star}(x)}^{-1}$ for Problem 3 (18) has the following properties: (i) it preserves the eigenvectors of $M(x)$, which define the principal axes of the associated confidence ellipsoid; (ii) it matches $M(x)$ along directions corresponding to eigenvalues within the range $\lambda_{\text{min}},\lambda_{\text{max}}]$; and (iii) it only adjusts the radii of the ellipsoid along the remaining principal axes to satisfy the constraint. This is visualized in Figure 2. Similarly, the confidence ellipsoid associated to (16) is obtained by projecting $M(x)$ onto the set of diagonal matrices (i.e., axis-aligned ellipsoids). Finally, in the case of (20), the radii of the axis-aligned ellipsoid are adjusted to satisfy the eigenvalue constraint.
@@ IV-C2 Preprocessed Measurements and Non-Additive Noise
In SLAM, raw measurements are often preprocessed and transformed nonlinearly into standard models supported by popular solvers. For instance, raw range-bearing measurements (corrupted by additive noise with covariance $\Sigma_{\text{true}}$) are often expressed in Cartesian coordinates. As a result, although the raw measurements generated by the sensor have identically distributed noise, the transformed measurements that appear in the least squares problem may have different covariances because of the nonlinear transformation. Let $\Sigma_{\text{true}}$ be the covariance of raw measurements, and $\Sigma_{i}$ be the covariance of the $i$th transformed measurement. In practice, $\Sigma_{i}$ is approximated by linearization, i.e., $\Sigma_{i}\approx J_{i}\Sigma_{\text{true}}J_{i}^{\top}$ in which $J_{i}$ is the (known) Jacobian of the transformation. It is easy to verify that Theorem 1 readily extends to this case when $J_{i}$’s are full-rank square matrices by replacing the sample covariance $S(x)$ as defined in (3) with
@@ V-B Block-Coordinate Descent
    Fix $P$ to its most recent value and minimize the joint MAP objective function in Problems 1-4 with respect to $x\in\mathcal{M}$. This results in a standard nonlinear least squares problem (where residuals are weighted by $P$) over $\mathcal{M}$, which can be (locally) solved using existing solvers such as [3, 13, 20] (see Remark 2).
@@
Two variants of this procedure are shown in Algorithms 3 and 4. In Step 1 of Algorithm 3, $R(x,P)\triangleq\langle M(x),P\rangle$ denotes the component of the joint cost function $F(x,P)$ that depends on $x$. As demonstrated in Remark 2, minimizing this function with respect to $x$ for a fixed $P$ is equivalent to minimizing the associated weighted nonlinear least squares objective. Moreover, while Algorithm 3 uses Riemannian gradient descent to update $x$ in Step 1, in principle one can use any (trust-region or line-search) optimization method that produces a descent iteration such as those already implemented in [3, 13, 20]. [^2] We analyze the convergence properties of Algorithm 3 in Theorem 3 when Riemannian gradient descent is used in Step 1.. Algorithm 4 is applicable to problems where in Step 1 of BCD one can exactly minimize $F(x,P)$ over $x$ and find the (unique) minimizer $x^{t}$ for the current value of $P$.
The BCD algorithms of the type considered here address the limitations of Algorithm 2. Specifically, the problem in the first step can be readily solved using standard solvers widely used in robotics and computer vision, which are also capable of exploiting sparsity in residuals. The second step is highly efficient, as it only requires computing the $m\times m$ matrix $M(x)$ as defined in (5), which can be done in $O(k\cdot m^{2})$ time. In the case of Problem 3, one must also compute the eigendecomposition of $M(x)$. In practice, $m$ (the dimension of the residuals) is typically a small constant (i.e., $m=O(1)$; e.g., in 3D PGO, $m=6$), and therefore the overall time complexity of the second step of BCD is $O(k)$, i.e., linear in the number of measurements. In a 2D PGO problem (i.e., $m=3$) with k=5,598k=5{,}598 measurements, updating the information matrix in Step 2 takes about one millisecond on a laptop CPU.
@@
[…]
4: // Initialize the information matrix
[…]
6:while $t\leq\tau$ do
[…]
8: // Step 1: Update $x^{t}$ using a descent step using retraction $\text{Retr}_{x^{t-1}}(\cdot)$, Riemannian gradient $\mathrm{grad}_{x}\,R(\cdot,\cdot)$ with respect to $x$, and step-size $\eta$ where $R(x,P)\triangleq\langle M(x),P\rangle$; see Theorem 3 and Appendix E.
[…]
10: // Step 2: optimize $P$ (Theorem 1)
[…]
12:end while
13:return $(x^{t},P^{t})$
14:end procedure
Algorithm 4 Block-Exact BCD for Joint MAP
1:procedure BCD($x_{\text{init}}$)
[…]
@@
[…]
5:while not converged do
[…]
7: // Step 1: optimize $x$
[…]
9: // Step 2: optimize $P$ (Theorem 1)
[…]
@@
Before studying the convergence properties of these algorithms we introduce the necessary assumption below. See Appendix E for background information.
@@
See Appendix F. ∎
As can be seen in Algorithm 4, one might be able to solve the optimization problems associated with each of the coordinate blocks uniquely and exactly. For example, for the case where the residuals $r_{i}(x)$ are affine functions of $x$ and $\mathcal{M}$ is convex, the optimization problem associated with $x$ has a unique solution (assuming a non-singular Hessian) and can be solved exactly. This case satisfies the following assumption.
@@
The proof of the theorem follows directly from [25, Theorem 1]. ∎
@@ VI-A Linear Measurement Model
in which $\Sigma^{\text{base}}\succeq 0$ is a fixed random covariance matrix, and $\sigma^{2}\in\{10^{-2},10^{-1},1,10,10^{2}\}$ is a variable that controls the noise level in our experiments. We did not impose a prior on $P$, leading to unconstrained joint ML estimation of $x_{\text{true}}$ and $\Sigma_{\text{true}}$.
@@ VI-B Pose-Graph Optimization Ablations
Dataset: We used a popular synthetic PGO benchmark, the Manhattan dataset [24], and generated new measurement realizations with varying values of the actual noise covariance matrix. The dataset consists of 3,5003{,}500 poses and k=5,598k=5{,}598 relative-pose measurements. This dataset is notoriously poorly connected [19]. Therefore, to analyze the effect of connectivity on covariance estimation, we also performed experiments on modified versions of this dataset, where additional loop closures were introduced by connecting pose $i$ to poses $i+2$ and $i+3$. This modification increased the total number of measurements to k=12,593k=12{,}593.
@@
Algorithms: We implemented a variant of Algorithm 3 in C++, based on g2o [20], for PGO problems. Instead of using Riemannian gradient descent, we used g2o’s implementation of Powell’s Dog-Leg method [26] on $\mathrm{SE}(2)$ to optimize $x$. In each outer iteration, our implementation retrieves the residuals for each measurement, computes $M(x)$ as defined in (5), and updates the noise covariance estimate using Theorem 1. We set $\lambda_{\text{min}}=10^{-4}$ and $\lambda_{\text{max}}=10^{4}$ in all experiments to handle cases where the smallest eigenvalue of $S(x)$ is (approximately) zero. We then perform a single iteration of Powell’s Dog-Leg method [26] to update the primary parameters $x$, initializing the solver at the latest estimate of $x$. To ensure convergence across all Monte Carlo trials, we ran 13 outer iterations. In this dataset, the per-iteration computational overhead of our algorithm (i.e., time spent updating the covariance matrix in Step 2 of Algorithm 3 relative to g2o with a given covariance) ranged from 0.9 to 1.5 milliseconds on an Intel i7-6820HQ CPU, which is negligible.
@@
    BCD: In line 11 of Algorithm 3, the noise information matrix is estimated using the ML estimate (i.e., no prior) subject to eigenvalue constraints. This is equivalent to (18) after replacing $M(x)$ with the sample covariance $S(x)$ at the current value for $x$.
@@
    BCD (diag): In line 11 of Algorithm 3, the noise information matrix is estimated using the ML estimate (as above), subject to both diagonal and eigenvalue constraints. This is equivalent to (20) after replacing $M(x)$ with the sample covariance $S(x)$ at the current value for $x$.
@@
    BCD (Wishart): In line 11 of Algorithm 3, the information matrix is updated according to (18); i.e., using the MAP estimate with a Wishart prior and under the eigenvalue constraints.
@@
    BCD (diag+Wishart): In line 11 of Algorithm 3, the noise information matrix is updated according to (20); i.e., using the MAP estimate with a Wishart prior and under the diagonal and eigenvalue constraints.
@@
The results indicate that, in most cases, diagonal BCD variants yield lower errors. This is expected, as the true noise covariance matrices are diagonal, and enforcing this prior information enhances estimation accuracy. Additionally, the MAP estimates obtained by BCD (Wishart) and BCD (diag+Wishart) outperform their ML-based counterparts. This is due to the fact that the eigenvalues of the sample covariance matrix $S(x)$ can be very small, indicating insufficient information for accurate noise covariance estimation. In such cases, the eigenvalue constraint effectively prevents a complete collapse of the estimated covariance. Interestingly, this issue was not encountered in [33] for ML estimation of the noise covariance matrix in PnnP. This discrepancy may suggest additional challenges in estimating noise covariance in SLAM (and related) problems, which typically involve a larger number of variables and sparse, graph-structured relative measurements. In such cases, incorporating a prior estimate may be necessary for accurate identification of noise covariance matrices. Finally, Figure 5(b) illustrates that the MAP estimation error is lower for the noise covariance of loop closures compared to odometry edges. This difference is partly because the prior value based on $\Sigma_{0}$ is closer to the true noise covariance matrix for loop closures.
@@ VI-C RIM Dataset
In this section, we present qualitative results on RIM, a large real-world 3D PGO dataset collected at Georgia Tech [11]. This dataset consists of 10,19510{,}195 poses and 29,74329{,}743 measurements. The g2o dataset includes a default noise covariance matrix, but with these default values, g2o struggles to solve the problem. Specifically, the Gauss-Newton and Dog-Leg solvers terminate after a single iteration, while the Levenberg-Marquardt solver completes 10 iterations but produces the trajectory estimate shown in Figure 6(a).
@@ VII Limitations
Algorithm 3, in its current form, is not robust to outliers, which limits its applicability in real-world scenarios. However, existing robust M-estimation techniques can be readily adapted for use with BCD. These methods iteratively reweight residuals based on a chosen robust cost function. Our preliminary results indicate that incorporating these weights into Step 1 of BCD (as is standard) and also into Step 2 (by replacing the sample covariance matrix $S(x)$ with a weighted average of residuals) yields an outlier-robust variant of BCD for the joint estimation problem. We are currently evaluating this approach and aim to implement it in real-time LiDAR and visual SLAM and odometry systems such as [10, 29].
Additionally, Theorem 3 requires $\mathcal{M}$ to be compact. This result therefore does not immediately apply to problems where $x$ contains translational components. However, we believe we can address this issue in SLAM and many other geometric estimation problems by confining translational components to potentially large but bounded subsets of $\mathbb{R}^{d}$ (where $d\in\{2,3\}$ is the problem dimension).
Finally, we acknowledge the extensive research on IGLS and Feasible GLS (FGLS) conducted as early as 1970s in econometrics and statistics, which remain largely unknown in engineering; see relevant references in [28, Chapter 12.5] and [33]. Exploring this rich literature may result in new insights and methods that can further improve noise covariance estimation in SLAM and computer vision applications.